# Region-Adaptive Transform with Segmentation Prior for Image Compression

### v2

@@ 4.2 Rate-Distortion Performance
We compare our SegPIC with previously well-performing methods, including discretized Gaussian mixture likelihoods and attention modules [12] (denoted as Cheng2020), ChARM [34] (denoted as Minnen2022), ChARM with WAM [49] (denoted as Zou2022), image-specific data-dependent transform model [43] (denoted as Wang2022), the best-performing traditional method VTM-17.0 [24] and influential traditional method BPG [7]. Furthermore, we introduce two simplified models without self-attention and auto-regressive mechanism, denoted as baseline-small (BL-S) and the one incorporating our proposed SAL and RAT (Ours-S), to demonstrate the generality of our modules. Details about the architecture of these two models will be provided in supplementary materials.
As illustrated in Fig. 6 and Tab. 1, compared to VTM-17.0, our SegPIC could save 8.18% BD-rate [8] on Kodak and 8.26% on CLIC, which outperforms all the LIC and traditional methods mentioned above. To further verify the efficiency of the proposed RAT and SAL, We also compare SegPIC with the baseline Zou2022, which can save 5.45% BD-rate on Kodak, and 8.50% on CLIC (see Tab. 2). Moreover, Ours-S even achieves closely matched performance compared to ChARM. The MS-SSIM results are provided in supplementary materials.

### v4

@@ 4.2 Rate-Distortion Performance
We compare our SegPIC with previously well-performing methods, including discretized Gaussian mixture likelihoods and attention modules [12] (denoted as Cheng2020), ChARM [34] (denoted as Minnen2022), ChARM with WAM [49] (denoted as Zou2022), image-specific data-dependent transform model [43] (denoted as Wang2022), the best-performing traditional method VTM-17.0 [24] and influential traditional method BPG [7]. Furthermore, we introduce two simplified models without self-attention and auto-regressive mechanism, denoted as baseline-small (BL-S) and the one incorporating our proposed SAL and RAT (Ours-S), to demonstrate the generality of our modules. Details about the architecture of these two models will be provided in the appendix.
As illustrated in Fig. 6 and Tab. 1, compared to VTM-17.0, our SegPIC could save 8.18% BD-rate [8] on Kodak and 8.26% on CLIC, which outperforms all the LIC and traditional methods mentioned above. To further verify the efficiency of the proposed RAT and SAL, We also compare SegPIC with the baseline Zou2022, which can save 5.45% BD-rate on Kodak, and 8.50% on CLIC (see Tab. 2). Moreover, Ours-S even achieves closely matched performance compared to ChARM. The MS-SSIM results are provided in the appendix.
@@ Acknowledgements
Appendix
6 Network Design of Ours-S and BL-S
The detailed designs of two simplified models Ours-S and BL-S are illustrated in Fig. 10. The Hyper Decoder consists of two modules, Hyper Decoder Scales and Hyper Decoder Means, which have identical network structures and each generates the estimated $\sigma$ and $\mu$ for $\hat{y}\sim\mathcal{N}(\mu,\sigma^{2})$. The RAT after the Hyper Decoder also consists of both the scale and mean portions, which is not fully depicted in Fig. 2 (a) and Fig. 10 for brevity. The network architectures of Hyper Encoder and Decoder are shown in Tab. 4.
Figure 10: The overall framework of Ours-S and BL-S. We construct these two simplified models without self-attention and auto-regressive mechanisms. FM is a Factorized Model [5]. The details of Hyper Encoder and Decoder are shown in Tab. 4.
7 Rate-Distortion Performance on MS-SSIM
We also compare the MS-SSIM results with other LIC methods and VTM-17.0. Here we additionally introduce some LIC methods, e.g. Factorize Model [5] (denoted as Balle2017), Hyperprior Entropy Model [6] (denoted as Balle2018), Joint Autoregressive and Hierarchical Priors Model [33] (denoted as Minnen2018). Fig. 11 shows that our SegPIC outperforms these methods.
Table 4: The network architecture of Hyper Encoder, Hyper Doecoder Scales, and Hyper Decoder Means. “Conv 256,3,2↓" means the convolution operation with out-channel 256, kernel size 3 and downsample step 2. “TConv" means transposed convolution.
Hyper Encoder Hyper Decoder Scales Hyper Decoder Means
Conv 320,3,1 Conv 192,3,1 Conv 192,3,1
GELU GELU GELU
Conv 288,3,1 TConv 224,3,2↑ TConv 224,3,2↑
GELU GELU GELU
Conv 256,3,2↓ Conv 256,3,1 Conv 256,3,1
GELU GELU GELU
Conv 224,3,1 TConv 288,3,2↑ TConv 288,3,2↑
GELU GELU GELU
Conv 192,3,2↓ Conv 320,3,1 Conv 320,3,1
Figure 11: The rate-distortion performance on Kodak. The metrics are bpp↓ and MS-SSIM↑. All the LIC methods are optimized by loss function Eq. 3, where the distortions can be formulated as $D=(1-\text{MS-SSIM}(x,\hat{x}))$.